What is the basic difference in the nature of
the electric fields produced by static charges
and those induced by changing magnetic
fields ?
1
Expert's answer
2019-05-06T10:34:05-0400

The only difference is in the structure of the fields:

  • The field contribution from static charges are non-rotational (curl-free) but divergent.
  • The field contributions from time-varying magnetic fields are divergence-less but have non-zero curl.

You can read this right out of the Gauss and Faraday's Laws in their differential form:


"\text{div} \vec{E} = \nabla \cdot \vec{E} = \frac{\rho}{\epsilon_0}, \qquad 
\text{rot} \vec{E} = \nabla \times \vec{E} = - \frac{\partial \vec{B}}{\partial t}"